# Conserved currents in five-dimensional proposals for lattice chiral gauge theories Maarten Goltermana and Yigal Shamirb

### v1

@@ I Introduction
Recently, another proposal was put forward by Kaplan [3] and by Kaplan and Sen [4]. In the new proposal the slab geometry is replaced by a disk geometry, and the rim of the disk is identified with one of the four physical dimensions. [^1] The remaining three physical dimensions are cartesian. While we focus on $4+1$ dimensions in this paper, the discussion generalizes straightforwardly to $2+1$ dimension. Remarkably, the rim supports a single Weyl fermion of one chirality only [3]. Moreover, this feature appears to survive lattice discretization [4], thereby circumventing the no-go theorems [9, 10]. By construction, the dynamical gauge field resides on the rim of the disk, and is extended into the whole disk via a radial gradient flow, once again maintaining gauge invariance.
@@ II Conserved currents in the Grabowska–Kaplan framework
These currents, too, are gauge invariant and conserved. The ${\rm U}(1)$ symmetry associated with the vector current ${\cal V}_{\mu}$ rotates $\psi_{R}$ and $\psi_{L}$ with the same phase. This is baryon number symmetry, a good symmetry of (one-flavor) QCD. The other $U(1)$ symmetry, associated with the current ${\cal A}_{\mu}$, rotates $\psi_{R}$ and $\psi_{L}$ with opposite phases; this is the axial symmetry, which is anomalous in QCD. The problem is thus that in the GK framework the axial symmetry is an exact symmetry, too. As a result, also the (singlet) axial current ${\cal A}_{\mu}$ is both conserved and gauge invariant within the GK framework. This generates a conflict between the standard properties of (one-flavor) QCD, and the features of its GK formulation. If the GK framework for regulating QCD leads to a local and Lorentz invariant continuum limit, that continuum limit must be different from the one obtained from any of the standard lattice regularizations of QCD.
@@
By contrast, within the GK formulation of the one-flavor theory, the gauge invariant current ${\cal A}_{\mu}$ has no anomaly, and, after taking the continuum limit, we obtain the WTI333 At finite lattice spacing, $p_{\mu}$ is replaced by $(2/a)\sin(ap_{\mu}/2)$.
@@
Note that now $\Sigma$ is a true order parameter, since the axial symmetry is exact in the GK formulation. Assuming that chiral symmetry breaking takes place and $\Sigma\neq 0$, the (euclidean space) correlator $\left\langle{\cal A}_{\mu}\,\eta\right\rangle$ must now have a massless pole in order to saturate this WTI. After the analytic continuation to Minkowski space, if the theory is local and Lorentz invariant, so that its two-point correlation functions admit a spectral representation, then the massless pole is recognized as a NGB. This new NGB, being a singlet pseudoscalar meson, would signal a breakdown of universality. This is our main conclusion.
@@
The existence of a singlet pseudoscalar NGB in the physical spectrum, which is absent from the standard formulation of QCD, is a completely general phenomenon within the GK formulation. Generalizing the lattice setup to QCD with $N_{f}$ flavors, it is easy to see that the global symmetry of the GK formulation will be ${\rm U}(N_{f})_{L}\times{\rm U}(N_{f})_{R}$, and not ${\rm SU}(N_{f})_{L}\times{\rm SU}(N_{f})_{R}\times{\rm U}(1)_{B}$ as expected. Once again, we will have a singlet axial current which is both gauge invariant and conserved, and not anomalous. Again, assuming that the theory confines and breaks its chiral symmetry, and that it has a consistent continuum limit, then, in addition to the expected pions for $N_{f}\geq 2$, there will be a superfluous, singlet pseudoscalar NGB.
@@ IV Dynamical considerations: the fate of ’t Hooft vertices
What is the dynamics responsible for the existence of such additional zero modes? We conjecture that one way for them to arise is as follows. As explained above, under the class of lattice flows we introduce in the appendix, the size of the instanton keeps shrinking with the fifth coordinate. After a long enough flow, the instanton’s size will become comparable to the lattice spacing $a$, which enables it to eventually disappear altogether. The 5D point $(x,s)$ inside the bulk where the instanton disappears must exhibit a dislocation of the flowed gauge field. We conjecture that a new zero mode may develop with support on this dislocation.888 For related setups where a bulk zero mode appears, see Refs. [15, 16]. In order to restore the conservation of the individual fermion numbers of the 5D fields, the bulk zero modes must have opposite U(1) charges from the 4D zero modes. Unlike the power-law decay of the familiar instanton zero modes, we expect the bulk zero modes to be exponentially localized.999 Such bulk zero modes would resemble the localized zero modes of the supercritical Wilson kernel of DWFs [17, 18, 19].
@@
for some constant $C$, which in turn is independent of $s$. Here $M=O(a^{-1})$ is the mass of the bulk 5D fermions.111111 Also known as the domain-wall height. The bound arises because the propagator in the fifth direction falls off like $e^{-M|s-s^{\prime}|}$, and the effective 4D fields are located on the boundary $s^{\prime}=0$. As long as the instanton’s size is large compared to the lattice scale,121212 See the appendix for details. one expects that the product $Ms$ diverges in the continuum limit. Hence, the expectation value of $O_{\rm tot}$ vanishes.
@@
In the above argument we have used that the contraction $\langle\Psi^{(L)}(x,s)\overline{\Psi}^{(R)}(x,s)\rangle$ vanishes identically, because it does not preserve the individual fermion numbers of the 5D fields. The same result generalizes to the thermodynamical limit. In order to take the thermodynamical limit we introduce a small mass term $m_{q}(\overline{\psi}_{R}\psi_{L}+\overline{\psi}_{L}\psi_{R})$. Since the mass term couples the two 5D fields, now $\langle\Psi^{(L)}(x,s)\overline{\Psi}^{(R)}(x,s)\rangle$ is non-zero, and (quark-disconnected) terms that include this contraction must be considered as well. However, any propagation from one 5D field to the other must go through the mass insertion, which in turn lives on the boundary. Hence131313 The  $\stackrel{{\scriptstyle\textstyle<}}{{\sim}}$  symbol indicates the presence of a proportionality factor, similarly to Eq. (13).
@@ V Discussion
When the target 4D theory is QCD with a small enough $N_{f}$ to enable chiral symmetry breaking, we concluded that the spectrum will contain a superfluous Nambu-Goldstone boson, which is a singlet pseudoscalar meson. Apart from chiral symmetry breaking, the further assumptions leading to this conclusions include Lorentz invariance and locality; all of them are well established properties of QCD.
@@
We comment that the so-called “Symmetric Mass Generation” (SMG) approach141414 For a review of SMG, and its relation to lattice chiral gauge theories, see Ref. [28]. to the construction of lattice chiral gauge theories does not have a similar U(1) issue. While gauge invariance is always maintained, the U(1) issue is avoided. The reason is that the multi-fermion and/or Yukawa interactions introduced into the lattice action are designed to break explicitly any symmetry not present in the target continuum theory, including the fermion number symmetry. Nevertheless this approach may fail for various dynamical reasons. Recently, we pointed out that the SMG approach has a potential issue with unwanted propagator zeros that could take the form of ghost states in the continuum theory. For details, as well as references to the original literature, see Ref. [29].
It is an open question whether or not the proposals of Refs. [2, 3] can lead to a consistent 4D quantum field theory in the continuum limit. If they do, this would signal a breakdown of universality, in the following sense. Considering once again the QCD example, the reason is that the new universality class would contain a singlet pseudoscalar Nambu–Goldstone boson in the physical spectrum if chiral symmetry breaking takes place, whereas the standard formulation of QCD does not have such a Nambu–Goldstone boson. If chiral symmetry breaking does not take place, and/or the 4D theory is not local (or Lorentz invariant), this would also mean that the theory is different from QCD as obtained in the continuum limit of one of the standard lattice regularizations. A similar breakdown of universality would happen in the case of a chiral gauge theory, because in the proposals of Refs. [2, 3] there exists a conserved and gauge invariant fermion number current, while Lüscher’s approach does not have such a current.
@@ Appendix A Lattice gradient flows
On the lattice, the action that generates the flow151515 For the lattice version of the flow equation, see Ref. [8]. can be expanded as a power series in the (squared) lattice spacing $a^{2}$, much like in the Symanzik improvement program [30, 31, 32, 33, 34]. Here we will limit ourselves to the two simplest lattice gauge actions. The first is the single plaquette (Wilson) action
@@
where $P_{x,\mu\nu}^{\rm rect}$ is the oriented product of links around the rectangle, with $\mu$ ($\nu$) corresponding to the short (long) side of the rectangle.161616 Equation (17) uses the normalization of the rectangle term most common in the literature. Expanding to $O(a^{2})$, these actions are given by [31, 32]

### v2

@@ I Introduction
Recently, another proposal was put forward by Kaplan [3] and by Kaplan and Sen [4]. In the new proposal the slab geometry is replaced by a disk geometry, and the rim of the disk is identified with one of the four physical dimensions. [^1] The remaining three physical dimensions are cartesian. While we focus on $4+1$ dimensions in this paper, the discussion generalizes straightforwardly to $2+1$ dimensions. Remarkably, the rim supports a single Weyl fermion of one chirality only [3]. Moreover, this feature appears to survive lattice discretization [4], thereby circumventing the no-go theorems [9, 10]. By construction, the dynamical gauge field resides on the rim of the disk, and is extended into the whole disk via a radial gradient flow, once again maintaining gauge invariance.
@@ II Conserved currents in the Grabowska–Kaplan framework
These currents, too, are gauge invariant and conserved. The ${\rm U}(1)$ symmetry associated with the vector current ${\cal V}_{\mu}$ rotates $\psi_{R}$ and $\psi_{L}$ with the same phase. This is baryon number symmetry ${\rm U}(1)_{B}$, a good symmetry of (one-flavor) QCD. The other $U(1)$ symmetry, associated with the current ${\cal A}_{\mu}$, rotates $\psi_{R}$ and $\psi_{L}$ with opposite phases; this is the axial symmetry ${\rm U}(1)_{A}$, which is anomalous in QCD. The problem is thus that in the GK framework the axial symmetry is an exact symmetry, too. As a result, also the (singlet) axial current ${\cal A}_{\mu}$ is both conserved and gauge invariant within the GK framework. This generates a conflict between the standard properties of (one-flavor) QCD, and the features of its GK formulation. If the GK framework for regulating QCD leads to a consistent continuum limit, that continuum limit must be different from the one obtained from any of the standard lattice regularizations of QCD.
@@
By contrast, within the GK formulation of the massless one-flavor theory, the gauge invariant current ${\cal A}_{\mu}$ has no anomaly, and, after taking the continuum limit, we obtain the WTI333 At finite lattice spacing, $p_{\mu}$ is replaced by $(2/a)\sin(ap_{\mu}/2)$.
@@
Note that now $\Sigma$ is a true order parameter, since the axial symmetry ${\rm U}(1)_{A}$ is exact in the GK formulation. The next step is to decompose the correlator in terms of invariant amplitudes, requiring translation and Lorentz invariance. [^4] After the analytic continuation to Minkowski space. Thanks to the simple form of the correlator, it depends on only a single invariant amplitude,
$\left\langle{\cal A}_{\mu}\,\eta\right\rangle=-ip_{\mu}F(p^{2})\ .$ (6)
Substituting this back into Eq. (5), the unique solution is $F(p^{2})=\Sigma/p^{2}$, or equivalently,
$\left\langle{\cal A}_{\mu}\,\eta\right\rangle=-i\Sigma\,\frac{p_{\mu}}{p^{2}}\ .$ (7)
Provided that chiral symmetry breaking takes place and $\Sigma\neq 0$, this result exhibits the pole of the Nambu–Goldstone boson. This new NGB, being a singlet pseudoscalar meson, would signal a breakdown of universality in QCD. This is our main conclusion.
@@
The existence of a singlet pseudoscalar NGB in the physical spectrum, which is absent from the standard formulation of QCD, is a completely general phenomenon within the GK formulation. Generalizing the lattice setup to QCD with $N_{f}$ flavors, it is easy to see that the global symmetry of the GK formulation will be ${\rm U}(N_{f})_{L}\times{\rm U}(N_{f})_{R}$, and not ${\rm SU}(N_{f})_{L}\times{\rm SU}(N_{f})_{R}\times{\rm U}(1)_{B}$ as expected. Once again, we will have a singlet axial current which is both gauge invariant and conserved, and not anomalous. Again, assuming that the theory confines and breaks its chiral symmetry, and requiring translation and Lorentz invariance, then, in addition to the expected pions for $N_{f}\geq 2$, there will be a superfluous, singlet pseudoscalar NGB.
@@ IV Dynamical considerations: the fate of ’t Hooft vertices
What is the dynamics responsible for the existence of such additional zero modes? We conjecture that one way for them to arise is as follows. As explained above, under the class of lattice flows we introduce in the appendix, the size of the instanton keeps shrinking with the fifth coordinate. After a long enough flow, the instanton’s size will become comparable to the lattice spacing $a$, which enables it to eventually disappear altogether. The 5D point $(x,s)$ inside the bulk where the instanton disappears must exhibit a dislocation of the flowed gauge field. We conjecture that a new zero mode may develop with support on this dislocation.999 For related setups where a bulk zero mode appears, see Refs. [15, 16]. In order to restore the conservation of the individual fermion numbers of the 5D fields, the bulk zero modes must have opposite U(1) charges from the 4D zero modes. Unlike the power-law decay of the familiar instanton zero modes, we expect the bulk zero modes to be exponentially localized.101010 Such bulk zero modes would resemble the localized zero modes of the supercritical Wilson kernel of DWFs [17, 18, 19].
@@
for some constant $C$, which in turn is independent of $s$. Here $M=O(a^{-1})$ is the mass of the bulk 5D fermions.121212 Also known as the domain-wall height. The bound arises because the propagator in the fifth direction falls off like $e^{-M|s-s^{\prime}|}$, and the effective 4D fields are located on the boundary $s^{\prime}=0$. As long as the instanton’s size is large compared to the lattice scale,131313 See the appendix for details. one expects that the product $Ms$ diverges in the continuum limit. Hence, the expectation value of $O_{\rm tot}$ vanishes.
@@
In the above argument we have used that the contraction $\langle\Psi^{(L)}(x,s)\overline{\Psi}^{(R)}(x,s)\rangle$ vanishes identically, because it does not preserve the individual fermion numbers of the 5D fields. The same result generalizes to the thermodynamical limit. In order to take the thermodynamical limit we introduce a small mass term $m_{q}(\overline{\psi}_{R}\psi_{L}+\overline{\psi}_{L}\psi_{R})$. Since the mass term couples the two 5D fields, now $\langle\Psi^{(L)}(x,s)\overline{\Psi}^{(R)}(x,s)\rangle$ is non-zero, and (quark-disconnected) terms that include this contraction must be considered as well. However, any propagation from one 5D field to the other must go through the mass insertion, which in turn lives on the boundary. Hence141414 The  $\stackrel{{\scriptstyle\textstyle<}}{{\sim}}$  symbol indicates the presence of a proportionality factor, similarly to Eq. (15).
@@ V Discussion
When the target 4D theory is QCD with a small enough $N_{f}$ to enable chiral symmetry breaking, we concluded that the spectrum will contain a superfluous Nambu-Goldstone boson, which is a singlet pseudoscalar meson. Apart from chiral symmetry breaking, we used only translation and Lorentz invariance.
@@
We comment that the so-called “Symmetric Mass Generation” (SMG) approach151515 For a review of SMG, and its relation to lattice chiral gauge theories, see Ref. [28]. to the construction of lattice chiral gauge theories does not have a similar U(1) issue. While gauge invariance is always maintained, the U(1) issue is avoided. The reason is that the multi-fermion and/or Yukawa interactions introduced into the lattice action are designed to break explicitly any symmetry not present in the target continuum theory, including the fermion number symmetry. Nevertheless this approach may fail for various dynamical reasons. Recently, we pointed out that the SMG approach has a potential issue with unwanted propagator zeros that could take the form of ghost states in the continuum theory. For details, as well as references to the original literature, see Ref. [29].
It is an open question whether or not the proposals of Refs. [2, 3] can lead to a consistent 4D quantum field theory in the continuum limit. If they do, this would signal a breakdown of universality, in the following sense. Considering once again the QCD example, the reason is that the new universality class would contain a singlet pseudoscalar Nambu–Goldstone boson in the physical spectrum if chiral symmetry breaking takes place, whereas the standard formulation of QCD does not have such a Nambu–Goldstone boson. If chiral symmetry breaking does not take place, and/or the 4D theory is not translation or Lorentz invariant, this would also mean that the theory is different from QCD as obtained in the continuum limit of one of the standard lattice regularizations. A similar breakdown of universality would happen in the case of a chiral gauge theory, because in the proposals of Refs. [2, 3] there exists a conserved and gauge invariant fermion number current, while Lüscher’s approach does not have such a current.
@@ Appendix A Lattice gradient flows
On the lattice, the action that generates the flow161616 For the lattice version of the flow equation, see Ref. [8]. can be expanded as a power series in the (squared) lattice spacing $a^{2}$, much like in the Symanzik improvement program [30, 31, 32, 33, 34]. Here we will limit ourselves to the two simplest lattice gauge actions. The first is the single plaquette (Wilson) action
@@
where $P_{x,\mu\nu}^{\rm rect}$ is the oriented product of links around the rectangle, with $\mu$ ($\nu$) corresponding to the short (long) side of the rectangle.171717 Equation (19) uses the normalization of the rectangle term most common in the literature. Expanding to $O(a^{2})$, these actions are given by [31, 32]